Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day05


Crew 102 Log Book for March 30, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: suboptimal (head cold, light stomachaches)

Time Departed/Returned from EVA: No EVA

Brief Narrative of Field Mission Results:

EVA Data/Interpretations:

Bacteria samples analysis ongoing

We had problems in dowloading GPS track data for offline analysis with dedicated software.

We are going to use the GPS applications available on some crew smart-phones.

Engineering/Hab Maintenance:

Electrical power outage. Problem provisionally fixed by Don Lusco.

Honey generator not working

Water Heater leaking problem manageable.

We are currently try to save propane (low level, 12%).

Plans for Tomorrow:

Second long EVA aimed at the collection of additional rock samples for the cyanobacteria study  and for some geological studies.

Analysis of Radiotelescope data after calibration.

Attempt to bring back grey water system in operation based on the procedure from Mission Support Chief Engineer Continue on current science projects (psychological, physiological, Hub Monitoring and Alarming, Hub Structural Corrosion)

Dismount and packaging of telescope in order to hand it over to Don Lusco as instructed by Peter Detterline

Support Requested:

Waiting for propane refill.

Don Lusco for final fixes on power generator.

Miscellaneous:


Commander's Report

Franco Carbognani Reporting

The day has been characterized by a severe power outage.

Just after lunch power was gone.

We checked the Kitty generator and we found stopped and pretty warm.

We checked the oil and we found it at the minimal level. We supposed the generator had stopped for this reason. We refilled a bit too much the oil going over the maximum level.

Trying to restart the generator it would not work.

In the meantime we contacted Joshua via cellular phones and we have been suggested to wait for Don Lusco.

While waiting for him we tried to start Honey, the spare generator. We were finally able to run it but no power was available on the Hab.

Don arrived in the late afternoon, fixed our oil overfill, substituted the diesel filter which

was supposed to be the main cause of the engine failure and the generator would restart.

While making checks in the Hub the generator stopped again.

We finally realized that the problem was on the motor cooling system with the cooling fan not rotating properly and very low level of cooling liquid.

We have provisionally fixed the cooling fan problem by adding an external fan.

Don should be able to come back tomorrow evening for a final fix.

Don confirmed us that the Honey generator run but do not provide electrical energy.

For what concern the propane refill, Don confirmed this has been ordered at the beginning of the week and it should be coming soon.

Because of the generator problem many of the activities planned for today had to be postponed but some science projects could be carried a bit forward.

Mood of the crew is good in spite of the generator annoyance.

John most probably had fiver last night. While trying to measure it this morning we found that the thermometer available in the Hab is not well calibrated since it was reporting unrealistically low temperatures. It should be replaced.

John seems now stable  and do not feel fever at the moment.


Angeliki and Kavya stomachaches are getting a bit better. Against those, a cure based on "gnocchi" has been tried with some beneficial effects.

Lara has performed a microscope analysis of few water samples. She could not spot the presence of living organisms but the presence of solid particles not well identifiable.

We will continue boiling and filtering our drinking water.




Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Kitty stopped running and has been provisionally repaired with the support of Don Lusco.

We tried to use Honey as a backup generator but we discovered it has been dead for some time.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.70

Propane Tank (Full/0.75/0.5/0.25/Empty): 12%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75


Notes/Comments:

On Mission Support Chief Engineer indications, both GreenHab termostats have been set to 50 degree F° in order to save propane.


Oil:

Kitty Oil Quantity (# of Quarts in storage): Was at lowest level, refilled (a bit too much indeed....) now 4/4

Honey Oil Quantity (# of Quarts in storage):

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 16

Trailer Potable Water Tank Level (inches from bottom): 32

Hab Potable Water Tank Level (inches from bottom): 30

Potable Water Meter Reading: 44202,7

In to GreenHab Meter Reading:


Notes/Comments:

Water heater leaking problem: Discussed with Don Lusco who should soon try to fix it permanently. Manageable problem at the moment.


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Gray water system not functional at the moment.

To be noted the fact that during electrical outage time the heaters of the Hub were on all the time driving the temperature up to 104F°. This imply that the thermostats are non functional during this time. Should we close the propane valve going to the heaters in case of another electrical outage?


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no):yes

Tire Status: good

Odometer at end of day: 123761.4

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: n

Oil Checked: n

Fuel Consumed:

Tire Status: fine

Comments:


ATV 2 (Spirit):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: fine

Comments: this ATV does not start


ATV 3 (Viking I):

Used: n

Oil Checked: n

Fuel Consumed:

Tire Status: fine

Comments:


ATV 4 (Viking II):

Used: n

Oil Checked: n

Fuel Consumed:

Tire Status: fine

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 70

Thermostat setting downstairs (Farenheit):66

Actual temperature downstairs (Farenheit): 68


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:


Health and Safety Officer's Report

Dr. John E. Deaton Reporting

One crew member has contracted a head cold and cough as mentioned in a previous report, and stated he had a mild fever (MDRS thermometer was not calibrated properly so cannot confirm) the evening of March 29, but it seems to have subsided.  Cough is still present but not to the same degree as before. Two doses of Theraflu were administered March 30, and one dose today.

Other crew have not reported anything out of the ordinary.


Chef's Report

Franco Carbognani Reporting

1a) Today was

- a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

Yes, a special meal was made to help some of the crew-members recover from stomach aches

2)  Foods served at today's lunch

Potato Gnocchi with tomato sauce, no special beverage

3) List any main meal foods not finished by the crew. 

Baking soda was added to leftover dough from making gnocchi, and used to make potato bread for dinner.

4) "Gnocchi al Pomodoro" Recipe

2 cups dried potato

1/2 cup dried spinach

Rehydrate 1 hour with 6 c. boiling water, stirring at intervals. Process with the mixer.

Let cool then add flour until get a bread dough like consistency.

Out of the dough roll cylinders of approx. 1 cm diameter and cut them into 2 cm pieces.

Roll them on a surface with flour.

Cook them in boiling water as normal pasta. They are ready when they start floating to the water surface

Sauce

4 tbsp dried onions.

1 cup tomato powder

Fry the onions with a 1 tbsp olive oil until they become brown. Add some water 

and the tomato powder. Stir and let it cook for 20 minutes, stirring from time to time. 

5)  Lessons learned

You can make decent gnocchi out of those dried ingredients!


Journalist Report

Kavya Manyapu Reporting

It has been a bit of a quiet day today for Crew 102….until the power system of the Hab broke down early afternoon. The crew has tried to start the main diesel generator and the backup gas generator but have not been successful so far as the systems might have broken down. Initially it was learned that the oil level in the generator was low, but even after the oil was replaced the generators did not work. The mission support had a backup crew member in orbit to come help the crew on Mars. The back-up crew member has landed on Mars and trying to help crew 102. Apart from fixing the power generator the commander who is the head chef for today cooked a really delicious meal for the crew. The crew enjoyed Hab made Italian Gnocchi! The physiology and psychology tests continue to take place. All the crew is still sitting in the dark waiting for the power to turn on. Let’s wait and see how long they survive.

8.30pm

Hurrah! Crew 102 have been lucky and got back the electricity after a marathon of 8.5 hours. It was just learned that the radiator fan belt was broken. And a back up mechanism has been used to make the generator work. But not sure how long this is going to stay. Hope the crew gets some dinner to eat this night. The crew is still fit and healthy but could not work on many of the experiments planned for today including EVAs.

Let’s wish them good luck for tonight and the rest of the days they need to spend on Mars!